1. Jim's mass is m = 62 kg. He is standing in a lift, with no motion relative to the lift.

a) The lift accelerates upwards at 2.0 m.s^-2. What is the magnitude of the force the floor of the lift exerts on Jim?
b) Later, the lift accelerates downwards at 2.0 m.s^-2. What is the magnitude of the force the floor of the lift exerts on Jim?
c) Just suppose that the lift were to accelerate downwards at 10.0 m.s^-2. What force would the floor of the lift exert on Jim?
1
Expert's answer
2020-03-05T09:37:38-0500

a)


"N_1=m(a+g)=62(2+9.8)=730\ N"

b)


"N_2=m(g-a)=62(9.8-2)=480\ N"

c)


"N_3=0\ N"

For "g\leq a" there is no force between floor and Jim. He would accelerate upwards relative lift.